OpenStudy (waheguru):

icolate s for thise=14(H)+8%(s) confused in how to solve this helppp

OpenStudy (anonymous):

can you change 8% into a decimal

OpenStudy (waheguru):

yea

OpenStudy (anonymous):

what is that 14(H)+8%(s)=?

OpenStudy (waheguru):

i have to take s out to it becomes s= ???

OpenStudy (anonymous):

now, i am saying that you question is unclear, i cannot tell what it is equal to

OpenStudy (waheguru):

oh it is equal to e

OpenStudy (waheguru):

e=14(H)+8%(s

OpenStudy (waheguru):

plz show steps

OpenStudy (anonymous):

okay simply subtract 14(h) from both sides you should get: e-14(h)=8%(s) Then lets convert 8% to a decimal we should get: e-14(h)=.08(s) Now divide both sides by .08 and you should get: (e-14(h))/(.08)=s

OpenStudy (waheguru):

whhy is the 8% with s dont we move it also??

OpenStudy (anonymous):

the are multipled toghether: 8% *s You see. Now to do mulitplication we use divsion

OpenStudy (waheguru):

ok thx

OpenStudy (anonymous):

anytime